a cockroach that is 1.6 inches long can move 80 inches per second. how many times your height could you move if you were a cockroach? my height is 5'1
Answer by stanbon(75887)

a cockroach that is 1.6 inches long can move 80 inches per second. how many times your height could you move if you were a cockroach? my height is 5'1
----
Use a proportion:
move/length = move/length
--
move/61" = 80/1.6"
---
move = 61(80/1.6)inches per second
---
move = 3050 inches per second
---
times your height = 3050/61 = 50
===================================
Cheers,
Stan H.